August 14th 2023

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • Shower
      • still need to start cosmics data, Kate is interested to help
    • HCAL
      • took cosmics, looks ok a few channels need to be check
      • could not work because of crate work
      • missing TDC might be due to noise - still investigating - will present Wednesday - pattern using pulser, looking in coincidence trigger
    • Hodoscope
      • Ralph could get into the hall last week
      • remote HV control was not working, power cycled the crate, power supply for NINO card it locked, need to check how to unlock it with Carlos
      • Hydra : some labeling added for HYDRA - some plots will not labeled -
  • Discussion WAPP trigger
    • Is trigger still set up ? (Mark) Most likely but need to be checked and timed in at the TS
    • Is trigger adequate ?
    • standard trigger with higher threshold seems good enough
    • Proton trigger in BigBite
      • need to check thresholds and efficiency
      • 2.5 to 2.8 GeV, need to look at simulation
      • 200 to 500 MeV energy deposit ?
      • postpone plan
      • firware, might need to update firmware
      • ALL measurement, requires SBS GEMs in data stream